OpenStudy (anonymous):

Hey guys, can I get some help? I've been staring at this problem for ages... Find x and y.

TheSmartOne (thesmartone):

Hey there. it seems like your question is missing some information :)

OpenStudy (anonymous):

Yeah hold on, I'm having internet struggles.

OpenStudy (anonymous):



OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

use pythagorean theorem

OpenStudy (anonymous):

\(y^2=x^2 + (9\sqrt{3}^2)\)

OpenStudy (anonymous):

Also, \[sin(60) = \frac{9\sqrt{3}}{y}\]

TheSmartOne (thesmartone):

Well it is a special right triangle, so it would be good to know this about a 30-60-90 triangle.